# Weakly Supervised 3D Object Detection via Multi-Level Visual Guidance

### v2

@@ Abstract
Weakly supervised 3D object detection aims to learn a 3D detector with lower annotation cost, e.g., 2D labels. Unlike prior work which still relies on few accurate 3D annotations, we propose a framework to study how to leverage constraints between 2D and 3D domains without requiring any 3D labels. Specifically, we employ visual data from three perspectives to establish connections between 2D and 3D domains. First, we design a feature-level constraint to align LiDAR and image features based on object-aware regions. Second, the output-level constraint is developed to enforce the overlap between 2D and projected 3D box estimations. Finally, the training-level constraint is utilized by producing accurate and consistent 3D pseudo-labels that align with the visual data. We conduct extensive experiments on the KITTI dataset to validate the effectiveness of the proposed three constraints. Without using any 3D labels, our method achieves favorable performance against state-of-the-art approaches and is competitive with the method that uses 500-frame 3D annotations. Code and models will be made publicly available at https://github.com/kuanchihhuang/VG-W3D.
@@
A key feature of autonomous systems is the ability to perceive and localize 3D objects accurately in the surrounding environments [10, 8, 9, 26, 4], enabling the agent (e.g., vehicles) to make informed decisions and navigate safely. However, acquiring annotated 3D labels for training 3D object detection models poses challenges due to its cost and complexity, particularly when compared with the speed of labeling 2D boxes on visual data (e.g., 3-16 times slower, as shown in [30]). Hence, weakly supervised learning for 3D object detection has emerged as a practical approach to address the annotation bottleneck.
@@ 3.3 Output-Level Visual Guidance
It is worth mentioning that any object detected with both 2D and 3D bounding boxes should exhibit a high overlap, as illustrated in Figure 4. This implies that in weakly-supervised learning scenarios where the 3D bounding box is unknown, we can utilize the ground truth 2D box to supervise the 3D boxes predicted by the 3D detector.
@@ 4.1 Experimental Setup
Datasets. We evaluate the proposed method on the KITTI 3D object detection [6] dataset, which comprises 7481 images for training and 7518 images for testing. Following the protocol established in [3], we divide the training samples into a training set (3712) and a validation set (3769) for conducting experiments for the validation set and ablation studies. For the nuScenes dataset, we leave the details and results in the supplementary material.

### v3

@@ Abstract
Weakly supervised 3D object detection aims to learn a 3D detector with lower annotation cost, e.g., 2D labels. Unlike prior work which still relies on few accurate 3D annotations, we propose a framework to study how to leverage constraints between 2D and 3D domains without requiring any 3D labels. Specifically, we employ visual data from three perspectives to establish connections between 2D and 3D domains. First, we design a feature-level constraint to align LiDAR and image features based on object-aware regions. Second, the output-level constraint is developed to enforce the overlap between 2D and projected 3D box estimations. Finally, the training-level constraint is utilized by producing accurate and consistent 3D pseudo-labels that align with the visual data. We conduct extensive experiments on the KITTI dataset to validate the effectiveness of the proposed three constraints. Without using any 3D labels, our method achieves favorable performance against state-of-the-art approaches and is competitive with the method that uses 500-frame 3D annotations. Code will be made publicly available at: https://github.com/kuanchihhuang/VG-W3D.
@@
A key feature of autonomous systems is the ability to perceive and localize 3D objects accurately in the surrounding environments [12, 10, 8, 28, 9, 4], enabling the agent (e.g., vehicles) to make informed decisions and navigate safely. However, acquiring annotated 3D labels for training 3D object detection models poses challenges due to its cost and complexity, particularly when compared with the speed of labeling 2D boxes on visual data (e.g., 3-16 times slower, as shown in [33]). Hence, weakly supervised learning for 3D object detection has emerged as a practical approach to address the annotation bottleneck.
@@ 3.3 Output-Level Visual Guidance
It is worth mentioning that any object detected with both 2D and 3D bounding boxes should exhibit a high overlap [31, 25], as illustrated in Figure 4. This implies that in weakly-supervised learning scenarios where the 3D bounding box is unknown, we can utilize the ground truth 2D box to supervise the 3D boxes predicted by the 3D detector.
@@ 4.1 Experimental Setup
Datasets. We evaluate the proposed method on the KITTI 3D object detection [6] dataset, which comprises 7481 images for training and 7518 images for testing. Following the protocol established in [3], we divide the training samples into a training set (3712) and a validation set (3769) for conducting experiments for the validation set and ablation studies.
@@ 5 Conclusions
Acknowledgements. This work was supported in part by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) contract number 140D0423C0074. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.